if you are going 80 miles per/hr and you have 20 miles to your destination how many mins will it take to get to your destination?
Answer by jim_thompson5910(35256)

Distance = Rate * Time

D = r*t

20 = 80*t

20/80 = t

1/4 = t

t = 1/4

It takes you 1/4 of an hour.

1/4 hour = (1/4)*(60) = 15 minutes

It will take you 15 minutes to get to your destination.


Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim